A $1,000 corporate bond has a maturity date 20 years from now and a coupon rate of 6 percent paid annually. Calculate the value of the bond if the required rate of return is a) 4% b) 6% c) 8%.

Strip the bond into an interest only bond and a face value only bond. That is, create a bond that consists only of the coupon interest payments and one that consists only of the face value. Calculate the value of each of these new bonds with required rates of return of a) 4% b) 6% c) 8%. Which bond has more interest rate risk? Which bond has more default risk? Now say that instead of annual coupon payments, the corporation makes payments every six months. Recalculate the value of your stripped bonds. Explain the results.
Answer #1

Annually Value of Bond Interest payment anually Rate of return@4% Interest=1000*6% Interest=60 Principle payment=1000 DiscounSTRIPS- Interest Payment anually Rate of return@4% Value of bond =Present value of Interest+Principle value of Principle ValuSemi anually Value of Bond Interest payment semi anually Rate of return@4% Interest=1000*6%*6/12 Interest=30 Principle paymenSTRIPS - Interest Payment Semi anually Rate of return@4% Value of bond =Present value of Interest+Principle value of Principl